# Event Trojan: Asynchronous Event-based Backdoor Attacks

### v1

@@ Abstract
As asynchronous event data is more frequently engaged in various vision tasks, the risk of backdoor attacks becomes more evident. However, research into the potential risk associated with backdoor attacks in asynchronous event data has been scarce, leaving related tasks vulnerable to potential threats. This paper has uncovered the possibility of directly poisoning event data streams by proposing Event Trojan framework, including two kinds of triggers, i.e., immutable and mutable triggers. Specifically, our two types of event triggers are based on a sequence of simulated event spikes, which can be easily incorporated into any event stream to initiate backdoor attacks. Additionally, for the mutable trigger, we design an adaptive learning mechanism to maximize its aggressiveness. To improve the stealthiness, we introduce a novel loss function that constrains the generated contents of mutable triggers, minimizing the difference between triggers and original events while maintaining effectiveness. Extensive experiments on public event datasets show the effectiveness of the proposed backdoor triggers. We hope that this paper can draw greater attention to the potential threats posed by backdoor attacks on event-based tasks. Our code is available at https://github.com/rfww/EventTrojan.
@@ 1 Introduction
Different from conventional images, asynchronous event data consists of a variety of asynchronous events, as illustrated by the point sets in Fig. 1. For compatibility with existing deep networks, event data needs to be converted into image-like representations to serve as the inputs of deep networks [20, 25, 12, 53, 15]. A typical solution is to adopt a representation module with different task-specific models for classification [12], recognition [55], segmentation [42], etc. Once the image-like representation is accessed, it can be directly injected with malicious triggers to initiate backdoor attacks for downstream models using existing image backdoor approaches [13, 10]. However, as the image-like representation module and the downstream task-specific models are usually tightly bounded [36], attackers cannot get this module to launch the attack. Furthermore, since the original event streams still retain their original contents, any backdoors introduced into the event representation become invalid if the representation is reconstructed from the unaltered event streams.
@@
In this paper, we propose Event Trojan, which injects triggers into the event data to enable backdoor attacks with high stealthiness and effective attack capability. An event, the basic unit of event data, consists of x-y coordinates, timestamps, and polarities. Therefore, a feasible solution is to craft the event trigger with multiple events based on predefined spatial coordinates, timestamps, and polarities, i.e., immutable trigger. Then, the attackers only need to inject this trigger into event streams to conduct backdoor attacks for victim models. Although this simple trigger can effectively impair the performance of numerous event-based victim models, fixed settings of the immutable trigger lead to limited generalization ability in various cases. This is due to neglecting the event distribution in the original data. So, we propose learning the trigger from the original event data to ensure the trigger has adaptive content capable of poisoning various event data, i.e., mutable trigger. Meanwhile, we design a novel loss function to optimize this trigger for high stealthiness and effectiveness. As displayed in Fig. 1, the mutable trigger (green points) shows a more realistic event form. Extensive experiments demonstrate that our proposed Event Trojan can easily inject triggers into the asynchronous event data and initiate effective backdoor attacks, even when defended by state-of-the-art defense methods.
@@ 2.1 Event data
Event data-based learning has gained increasing focus due to the advantages of the bio-inspired sensor, the event camera, which captures moving objects with high dynamic range and temporal resolution, low time latency and power consumption [41, 12, 1, 37, 55]. Gehrig et al. [12] design a grid-based representation that transfers the event data stream to image-like representations, enabling many state-of-the-art vision models that can be easily worked on the event streams. Schaefer et al. [37] propose an asynchronous event-based graph neural network, which treats the events as temporally evolving graphs to avoid sacrificing the sparsity and high temporal resolution. Sun et al. [43] propose an unsupervised domain adaption method for semantic segmentation on event data, which motivates the segmentor to learn semantic information from labeled images to unlabeled events. In addition, event-based studies achieve satisfactory performances in image deblurring [16, 52], optical flow [39, 53], object recognition [18, 55], video reconstruction [35], stereo matching [50], etc. Although event-based methods have drawn more attention from researchers recently, limited security studies on this topic have been conducted.
@@ 2.2 Backdoor attack
The backdoor attack is a typical topic to study the vulnerability of deep models [13], which is very different from the adversarial attack in two terms: attack fuse and attack process [32]. Backdoor attacks, injecting a trigger into data samples to mislead the model outputting an attacker’s desired label, have been extensively studied for the model security of 2D-image models [48], 3D point cloud networks [21], neural radiance fields [7], natural language processing networks [38], speech recognizers [19], etc. Gu et al. [13] first studied the backdoor attack in the deep learning area, injecting a checkerboard pattern as the trigger to mislead the classifier to output a given label on the triggered data. Subsequently, Chen et al. [4] propose a physical instance-based backdoor attack model, which employs the daily used products as triggers to avoid human censorship. Besides, some people make efforts to explore novel trigger patterns to improve the stealthiness, such as object reflection [26], image structure [29, 49] and frequency perturbations [23, 10], etc. Apart from the classification task, the backdoor attack is also studied in terms of semantic segmentation [10], object detection [3], video recognition [51], facial recognition [47], etc. Although a wide range of backdoor attack methods have been proposed to examine security issues across various tasks, it is still impossible to directly poison asynchronous event data by existing methods to execute backdoor attacks. This is due to the asynchronous property of event data, which only records information about pixels with brightness changes exceeding a certain threshold.
@@ 3.1 Backdoor attack
Given a dataset ${\mathcal{D}}=\{d_{i},l_{i}\}_{i=1}^{N}$, where $d_{i}$ and $l_{i}$ indicate the input data and the corresponding label. The backdoor attack aims to learn a mapping function: $f_{\theta}(d_{i})\rightarrow l_{i}$ while changing this mapping to $f_{\theta}(d_{i})\rightarrow c$ if $d_{i}$ contains a trigger injected by ${T}(d_{i})$. $f_{\theta}(\cdot)$ is a deep model with its learnable parameters $\theta$. $c$ is the attacker-chosen label, which is employed to evaluate the attack effectiveness [6]. For training a backdoor model, attackers first need to poison some input data with a poison ratio $\rho$ and then train the model with both benign and poisoned samples. Ultimately, this model can output accurate predictions on the benign samples while giving malicious outputs (e.g., $c$) when the attacker injects the designed triggers into input data [23].
@@ 4 Methodology
where $F_{\{\theta,\omega\}}$ denotes the event vision model with learned parameters, $T(\cdot)$ indicates the trigger injection function. $c$ is the attacker-chosen label. We first discuss the threat model to event-based tasks and then introduce immutable and mutable triggers in the following sections to reveal the possible approaches for backdoor attacks on event-based vision models
@@ 4.2 Immutable trigger
The essence of the immutable trigger lies in the deliberate placement of malicious events at particular spatial locations and time stamps within various event data streams. After injecting these events, they can maintain consistent spatial positions and time stamps across different event streams (see Fig. 1 (d)). Given that the spatial dimension dictates the shape and the temporal dimension influences the pixel values in the event representation, these injected malicious events manifest as identical patterns in the representations. As shown in the $4^{th}$ column of Fig. 3, our immutable trigger still can show identical patterns across different asynchronous event data. We synthesize the immutable trigger manually with $(x,y)$ coordinates sampled from a predefined region, timestamp by $\alpha$, and polarity by $\beta$. Then, we inject it into the original event stream and modify the label as the attacker-desired target. Based on this strategy, we can generate more triggered samples according to the poison ratio $\rho$ to train a victim model.
@@ 4.3 Mutable trigger
As shown in the last column of Fig. 3, the mutable trigger possesses two characteristics: ❶ The malicious events inserted across various asynchronous event data streams possess identical spatial values, ensuring that the trigger patterns maintain the same shapes within the image-like representation; ❷ The inserted events are given adaptive time stamps (see the event trigger in Fig. 1, leading to trigger patterns with unique pixel values in the image-like representation. The entire methodology for embedding malicious triggers with adaptive time stamps is depicted in Algorithm 1. In this process, several events are strategically placed at time stamps deduced by a malicious events injector, referred to as $T_{\xi}(\cdot)$. First, we randomly sample $m$ time stamps from the original event as the input of $T_{\xi}(\cdot)$, which can generate a trigger $\mathcal{T}$ using the predicted malicious time stamps. Then, we inject this trigger into the original event and modify the corresponding labels according to the attacker’s targets. Finally, we train the deep classifiers and our trigger injector $T_{\xi}(\cdot)$ jointly to encourage the produced triggers that best suit the classifiers. This scheme can utilize the classifier to guide the injector $T_{\xi}(\cdot)$ to learn the adaptive triggers for different events. We design a trigger optimization loss function to ensure that $T_{\xi}(\cdot)$ can learn the unique characteristics along the assigned dimensions. Note that our trigger generator is optimized with the classifier only during the training phase; this does not imply that the generator is tied to the classifier during inference.
@@
where $\mathbf{t}=\{t_{i}\}_{i=1}^{m}$, indicating the sampled $m$ time stamps from the benign event data. $T_{\xi}(\mathbf{t})$ denotes the generated malicious time stamps. $\psi(\cdot)$ involves calculating the square difference between the malicious time stamps and clean time stamps in terms of the expectation and variance, respectively. The detailed training process is shown in line $8\sim 14$ of Algorithm 1.
@@ 4.4 Implementation details
Table 1: Quantitative comparison results of different triggers imposed on event data-based deep models. We show the R. triggers (Representation triggers) obtained by BadNets [13] and FIBA [10] in Fig. 3.
@@ 5.1 Setup
To validate the effectiveness of our methods, We use the N-Caltech101 dataset [30] and the N-Cars dataset [41] in our evaluation. N-Caltech101 [30] is an event-based version of the frame-based Caltech101 dataset [9], which is obtained by affixing the ATIS sensor [34] to a motorized pan-tilt unit to record the moved Caltech101 examples. N-Caltech101 [30] consists of 4,356 training samples, 2,612 validating samples, and 1,741 testing samples in 101 classes. The amount of data varies greatly among different categories. N-Cars (Neuromorphic-Cars) [41] is a real-world event dataset for recognizing whether a car is present in a scene. It is recorded using an ATIS camera [34] that is mounted on a car. According to the partition in [37], the N-Cars dataset [41] includes 8392 training samples, 2462 validation samples, and 8608 testing samples in two classes.
@@
For evaluating the effectiveness of the proposed methods comprehensively, we quantify the results of 22 popular classifiers with different network architectures (including ResNet [14], VGG [40], EfficientNet [45], Inception [44], ViT [8], Swin Transformer [27], DeiT [46]. All models are implemented by the official codes, with modifications made only to the input and output channels. EST [12] is selected as the event representation network.
@@ 5.3 Ablation studies
BadNets [13] FIBA [10] Immutable T. Mutable T.

### v2

@@ Abstract
As asynchronous event data is more frequently engaged in various vision tasks, the risk of backdoor attacks becomes more evident. However, research into the potential risk associated with backdoor attacks in asynchronous event data has been scarce, leaving related tasks vulnerable to potential threats. This paper has uncovered the possibility of directly poisoning the event data stream by proposing Event Trojan framework with two kinds of triggers, i.e., immutable and mutable triggers. Specifically, our two types of event triggers are based on a sequence of simulated event spikes, which can be easily incorporated into any event stream to initiate backdoor attacks. Additionally, for the mutable trigger, we design an adaptive learning mechanism to maximize its aggressiveness. To improve the stealthiness, we introduce a novel loss function that constrains the generated contents of mutable triggers, minimizing the difference between triggers and original events while maintaining effectiveness. Extensive experiments on public event datasets show the effectiveness of the proposed backdoor triggers. We hope that this paper can draw greater attention to the potential threats posed by backdoor attacks on event-based tasks.
@@ 1 Introduction
Different from conventional images, asynchronous event data consists of a variety of asynchronous events, as illustrated by the point sets in Fig. 1. For compatibility with existing deep networks, event data needs to be converted into image-like representations to serve as the inputs of deep networks [20, 25, 12, 53, 15]. A typical solution is to adopt a representation module with different task-specific models for classification [12], recognition [55], segmentation [42], etc. Once the image-like representation is accessed, it can be directly injected with malicious triggers to initiate backdoor attacks for downstream models using existing image backdoor approaches [13, 10]. However, as the image-like representation module and the downstream task-specific models are usually tightly bounded [36], attackers cannot get this module to launch the attack. Furthermore, since the original event stream still retains its original contents, any backdoors introduced into the event representation become invalid if the representation is reconstructed from the unaltered event stream.
@@
In this paper, we propose Event Trojan, which injects triggers into the event data to enable backdoor attacks with high stealthiness and effective attack capability. An event, the basic unit of event data, consists of x-y coordinates, timestamps, and polarities. Therefore, a feasible solution is to craft the event trigger with multiple events based on predefined spatial coordinates, timestamps, and polarities, i.e., immutable trigger. Then, the attackers only need to inject this trigger into event stream to conduct backdoor attacks for victim models. Although this simple trigger can effectively impair the performance of numerous event-based victim models, fixed settings of the immutable trigger lead to limited generalization ability in various cases. This is due to neglecting the event distribution in the original data. So, we propose learning the trigger from the original event data to ensure the trigger has adaptive content capable of poisoning various event data, i.e., mutable trigger. Meanwhile, we design a novel loss function to optimize this trigger for high stealthiness and effectiveness. As displayed in Fig. 1, the mutable trigger (green points) shows a more realistic event form. Extensive experiments demonstrate that our proposed Event Trojan can easily inject triggers into the asynchronous event data and initiate effective backdoor attacks, even when defended by state-of-the-art defense methods.
@@ 2.1 Event data
Event data-based vision has gained increasing focus due to the advantages of the bio-inspired sensor, the event camera, which captures moving objects with high dynamic range and temporal resolution, low time latency and power consumption [41, 12, 1, 37, 55]. Gehrig et al. [12] design a grid-based representation that transfers the event data stream to image-like representations, enabling many state-of-the-art vision models that can be easily worked on the event stream. Schaefer et al. [37] propose an asynchronous event-based graph neural network, which treats the events as temporally evolving graphs to avoid sacrificing the sparsity and high temporal resolution. Sun et al. [43] propose an unsupervised domain adaption method for semantic segmentation on event data, which motivates the segmentor to learn semantic information from labeled images to unlabeled events. In addition, event-based studies achieve satisfactory performances in image deblurring [16, 52], optical flow [39, 53], object recognition [18, 55], video reconstruction [35], stereo matching [50], etc. Although event-based methods have drawn more attention from researchers recently, limited security studies on this topic have been conducted.
@@ 2.2 Backdoor attack
The backdoor attack is a typical topic to study the vulnerability of deep models [13], which is very different from the adversarial attack in two terms: attack fuse and attack process [32]. Backdoor attacks, injecting a trigger into data samples to mislead the model outputting an attacker’s desired label, have been extensively studied for the model security of 2D-image models [48], 3D point cloud networks [21], neural radiance fields [7], natural language processing networks [38], speech recognizers [19], etc. Gu et al. [13] first study the backdoor attack in the deep learning area, injecting a checkerboard pattern as the trigger to mislead the classifier to output a given label on the triggered data. Subsequently, Chen et al. [4] propose a physical instance-based backdoor attack model, which employs the daily used products as triggers to avoid human censorship. Besides, some people make efforts to explore novel trigger patterns to improve the stealthiness, such as object reflection [26], image structure [29, 49] and frequency perturbations [23, 10], etc. Apart from the classification task, the backdoor attack is also studied in terms of semantic segmentation [10], object detection [3], video recognition [51], facial recognition [47], etc. Although a wide range of backdoor attack methods have been proposed to examine security issues across various tasks, it is still impossible to directly poison asynchronous event data by existing methods to execute backdoor attacks. This is due to the asynchronous property of event data, which only records information about pixels with brightness changes exceeding a certain threshold.
@@ 3.1 Backdoor attack
Given a dataset ${\mathcal{D}}=\{d_{i},l_{i}\}_{i=1}^{N}$, where $d_{i}$ and $l_{i}$ indicate the input data and the corresponding label. The backdoor attack aims to learn a mapping function: $f_{\theta}(d_{i})\rightarrow l_{i}$ while changing this mapping to $f_{\theta}(d_{i})\rightarrow c$ if $d_{i}$ contains a trigger injected by ${T}(d_{i})$. $f_{\theta}(\cdot)$ is a deep model with its learnable parameters $\theta$. $c$ is the attacker-chosen label, which is employed to evaluate the attack effectiveness [6]. For training a backdoor model, attackers first need to poison some input data with a poison ratio $\rho$ and then train the model with both benign and poisoned samples. Ultimately, this model can output accurate predictions on the benign samples while giving malicious outputs (e.g., the attacker-desired label $c$) when the attacker injects the designed triggers into input data [23].
@@ 4 Methodology
where $F_{\{\theta,\omega\}}$ denotes the event vision model with learned parameters, $T(\cdot)$ indicates the trigger injection function and $c$ represents the attacker-chosen label. We first discuss the threat model to event-based tasks and then introduce immutable and mutable triggers in the following sections to reveal the possible approaches for backdoor attacks on event-based vision models.
@@ 4.2 Immutable trigger
The essence of the immutable trigger lies in the deliberate placement of malicious events at particular spatial locations and time stamps within various event data streams. After injecting these events, they can maintain consistent spatial positions and time stamps across different event streams (see Fig. 1 (d)). Given that the spatial dimension dictates the shape and the temporal dimension influences the pixel values in the event representation, these injected malicious events manifest as identical patterns in the representations. As shown in the $4^{th}$ column of Fig. 3, our immutable trigger can still show identical patterns across different asynchronous event data. We synthesize the immutable trigger manually with $(x,y)$ coordinates sampled from a predefined region, timestamp by $\alpha$, and polarity by $\beta$. Then, we inject it into the original event stream and modify the label as the attacker-desired target. Based on this strategy, we can generate more triggered samples according to the poison ratio $\rho$ to train a victim model.
@@ 4.3 Mutable trigger
As shown in the last column of Fig. 3, the mutable trigger possesses two characteristics: ❶ The malicious events inserted across various asynchronous event data streams possess identical spatial values, ensuring that the trigger patterns maintain the same shapes within the image-like representation; ❷ The inserted events are given adaptive time stamps (see the event trigger in Fig. 1), leading to trigger patterns with unique pixel values in the image-like representation. The entire methodology for embedding malicious triggers with adaptive time stamps is depicted in Algorithm 1. In this process, several events are strategically placed at time stamps deduced by a malicious events injector, referred to as $T_{\xi}(\cdot)$. First, we randomly sample $m$ time stamps from the original event as the input of $T_{\xi}(\cdot)$, which can generate a trigger $\mathcal{T}$ using the predicted malicious time stamps. Then, we inject this trigger into the original event and modify the corresponding labels according to the attacker’s targets. Finally, we train the deep classifiers and our trigger injector $T_{\xi}(\cdot)$ jointly to encourage the produced triggers that best suit the classifiers. This scheme can utilize the classifier to guide the injector $T_{\xi}(\cdot)$ to learn the adaptive triggers for different events. We design a trigger optimization loss function to ensure that $T_{\xi}(\cdot)$ can learn the unique characteristics along the assigned dimensions. Note that our trigger generator is optimized with the classifier only during the training phase; this does not imply that the generator is tied to the classifier during inference.
@@
where $\mathbf{t}=\{t_{i}\}_{i=1}^{m}$ indicates the sampled $m$ time stamps from the benign event data. $T_{\xi}(\mathbf{t})$ denotes the generated malicious time stamps. $\psi(\cdot)$ involves calculating the square difference between the malicious time stamps and clean time stamps in terms of the expectation and variance, respectively. The detailed training process is shown in line $8\sim 14$ of Algorithm 1.
@@ 4.4 Implementation details
Table 1: Quantitative comparison results of different triggers imposed on event data-based deep models. We show the Representation triggers (R. triggers) obtained by BadNets [13] and FIBA [10] in Fig. 3.
@@ 5.1 Setup
To validate the effectiveness of our methods, we use the N-Caltech101 dataset [30] and the N-Cars dataset [41] in our evaluation. N-Caltech101 [30] is an event-based version of the frame-based Caltech101 dataset [9], which is obtained by affixing the ATIS sensor [34] to a motorized pan-tilt unit to record the moved Caltech101 examples. N-Caltech101 [30] consists of 4356 training samples, 2612 validating samples, and 1,741 testing samples in 101 classes. The amount of data varies greatly among different categories. N-Cars (Neuromorphic-Cars) [41] is a real-world event dataset for recognizing whether a car is present in a scene. It is recorded using an ATIS camera [34] that is mounted on a car. According to the partition in [37], the N-Cars dataset [41] includes 8392 training samples, 2462 validation samples, and 8608 testing samples in two classes.
@@
For evaluating the effectiveness of the proposed methods comprehensively, we quantify the results of 22 popular classifiers with different network architectures, including ResNet [14], VGG [40], EfficientNet [45], Inception [44], ViT [8], Swin Transformer [27], DeiT [46]. All models are implemented by the official codes, with modifications made only to the input and output channels. EST [12] is selected as the event representation network.
@@ 5.3 Ablation studies
BadNets FIBA Immutable T. Mutable T.